Q. A particle starts form rest on a straight line. First it accelerates at 4 m/s 2 for 10 seconds. Then for 10 seconds it moves with the acquired constant velocity. Then it decelerates at 8 m/s 2 to rest. Total distance covered during the entire journey is Choose one: 500 m 600 m 700 m 800 m

Q.
A particle starts form rest on a straight line. First it accelerates at 4 m/s2 for 10 seconds. Then for 10 seconds it moves with the acquired constant velocity. Then it decelerates at 8 m/s2 to rest. Total distance covered during the entire journey is
500 m
600 m
700 m
800 m
 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
For 1st phase from
V1 = u + a* t
V1 = 40 m/sec
V²1 = u² + 2 *a*s
S = 1600/8 = 200 m
 
Now for second phase
V2 = v1
Now S = V2 *t = 40 * 10 = 400 m
 
Gor third phase 
Initial velocity = V2 = 40 m
From V²3 = V²2 - 2*a*s
S = 1600/16 = 100 m
 
Hence total distance covered = 200 + 400 + 100 = 700 m
 
Regards
Arun (askIITians forum expert)